body weight 54 kg volume 20 dm of which material the body is made


1
Expert's answer
2021-09-21T06:33:49-0400

Converting volume to "m^3", obtain:


"V = 20dm^3 = 20\times 10^{-3}m^3=0.02m^3"


Thus, the density of the body is:


"\rho = \dfrac{m}{V} = \dfrac{54kg}{0.02m^3} = 2700\dfrac{kg}{m^3}"

Thus, the material might be aluminum.


Answer. Aluminum.